OpenStudy (anonymous):

solve by completing the square x^2+4x-21=0

OpenStudy (anonymous):

\[x^2+4x-21=0\]\[x^2+4x=21\]\[x^2+4x+4=21+4\]\[(x+2)^2=25\]\[x+2=\pm5\]\[x=-3\]or\[x=7\]